492×28? EXPlain pls?




Answer by MathLover1(20849)


492×28=.........first multiply by tents (in your case 2)


296 x 8 = ?
MULTIPLY by TENS: 2x2 = 4
WRITE AND CARRY: Write the "4" and carry the "0"
MULTIPLY TENS: 2 x 9 = 18
ADD: Add the carried amount. 18 + 0 = 18
WRITE AND CARRY: Write the "8" and carry the "1"
MULTIPLY THE HUNDREDS: 2 x 4 = 8
ADD: Add the carried amount. 8 + 1 = 9
WRITE: Write down the "9"
Answer: 492 x 2 = 984
now do same thing multiplying by ones, and move last digit one space to the right of 984 like this
984
----...on last dash you will put ones
492 x 28 = ?
MULTIPLY by ONES: 2x8 = 16
WRITE AND CARRY: Write the "6" and carry the "1"
MULTIPLY TENS: 9 x 8 = 72
ADD: Add the carried amount. 72 + 1 = 73
WRITE AND CARRY: Write the "3" and carry the "7"
MULTIPLY THE HUNDREDS: 4 x 8 = 32
ADD: Add the carried amount.32 +7 = 39
WRITE: Write down the "39"
Answer: 492 x 8 = 3936
now
984
----...on last dash you will put ones, and you have
984
3936...and add
__________
13776